The probabilities that a batch of computers will contain XX defective computers are are given below. Find the standard deviation for the probability distribution. Round answer to the nearest hundredth.

XP(X)

--------------------------------------------------------------

0|0.5333330

1|0.2333331

2|0.2333332

3|0.0000003

What is the standard deviation for this probability distribution?

SELECT ALL APPLICABLE CHOICES

A)σ=0.82260

B)σ=1.1235

C)σ=1.2243

D)σ=0.92135

E)σ=1.2134

F) None of These
Answer #1

Mean calculation

\bar{x} = \sum (x*P(x)) \\ = (0*0.5333330)+(1*0.2333331)+(2*0.2333332)+(3*0.0000003) \\ = 0+ 0.2333331+ 0.4666664+ 0.0000009 \\ = 0.7000004

standard deviation calculation

\sigma = \sqrt{\sum (x^2*P(x))-\bar{x}^2} \\ = \sqrt{[(0^2*0.5333330)+(1^2*0.2333331)+(2^2*0.2333332)+(3^2*0.0000003)]-(0.7000004 )^2} \\ = \sqrt{1.1666686 -0.7000004 ^2} \\ = 0.82260option A is correct